# LLMs, You Can Evaluate It! Design of Multi-perspective Evaluation for Reports in Security Operation Centers

### v2

@@ 1 Introduction
However, writing analysis reports for security alerts forces a heavy workload on SOC practitioners, because these reports must contain actionable insights [kokulu2019MatchedandMismatchedSOCs]. To reduce this workload, tools, such as large language models (LLMs), to introduce actionable insights into analysis reports are in high demand [albanese2025towardsAI-driven]. Although there are several tools to generate analysis reports for security alerts [loumachi2025advancing, michelet2024chatgpt], automatically generated reports often contain inaccurate information, which may degrade performance in SOCs [kokulu2019MatchedandMismatchedSOCs]. Even when the state-of-the-art LLMs are utilized, they often cause hallucinations such as inaccurate information [huang2025hallucination]. As another motivation, writing analysis reports to the exact requirements of report-writing guidelines is stressful [nepal2024burnout]. The reason is that evaluation of analysis reports should identify any lack of descriptions as well as judgments from the perspectives of veteran SOC practitioners [RYAN2012Quantifying]. Although the most common mitigation approach for the above stress is to constantly evaluate analysis reports [kokulu2019MatchedandMismatchedSOCs], evaluations of the reports need evaluation criteria based on expert knowledge [kerstenfield2025tier1analyst, kersten2023give]. Moreover, the current guidelines are often abstract and may fail to capture the evaluation criteria from the perspective of veteran SOC practitioners [johnson2016guide, agyepong2020towards]. Namely, even evaluations of analysis reports using LLMs are challenging due to a lack of enough evaluation criteria.
@@
We first design the Analyst-wise Checklist, which contains the evaluation criteria in its specific items to reflect SOC practitioners’ knowledge for evaluating analysis reports in the real world, as the answer to RQ1. To design this, we also conducted a literature review of 13 public guidelines and handbooks in the IT and OT domains and then semi-structured interviews with 15 SOC practitioners. We also conducted a validation test to confirm the checklist with multiple SOC practitioners. (See Section 3 for details.)
Next, based on the above checklist, we propose a novel framework, called Multi-perspective Evaluation System for Security Analysis using Llm Assistance (MESSALA), to maximize the evaluation of analysis reports in SOCs. In a nutshell, MESSALA guides LLMs in evaluating analysis reports by leveraging the Analyst-wise Checklist from multiple perspectives. As described in Section 4, MESSALA can imitate SOC practitioners’ cognition [zhong2016automate] to obtain expert knowledge from the checklist, and therefore, can return appropriate evaluation results with actionable feedback for each analysis report. When we conduct quantitative evaluations with MESSALA, we identify that it can rate analysis reports on the evaluation criteria compared with the existing methods [fu2023gptscore, liu2023g-eval]. Thus, to answer RQ2, we confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports from the perspective of veteran SOC practitioners. (See Section 5 for details.)
Third, we conduct qualitative evaluations to identify whether the feedback generated by MESSALA aligns with judgments from the perspective of veteran SOC practitioners. As a result, we confirm if MESSALA can provide more actionable feedback with both novice and veteran SOC practitioners than the existing methods, because it is more specific and easier to understand. It means that MESSALA also enables LLMs to qualitatively evaluate analysis reports through its feedback. (See Section 6 for details.)
@@
  • We design the Analyst-wise Checklist that contains evaluation criteria for evaluating analysis reports to reflect real-world SOC practitioners’ knowledge through literature review and interviews with SOC practitioners as the answer to RQ1.
  • We propose MESSALA, a novel framework that maximizes evaluation for each criterion and provides feedback from the perspective of SOC practitioners, and confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports under the evaluation criteria. This is identical to the answer to RQ2.
@@ 2 Related Work and Our Problem Setting
In this section, we describe related works in terms of LLMs and SOCs as background. We then describe evaluation of analysis reports in SOCs as the main problem setting.
@@ 2.0.1 Large Language Models
LLMs generate text outputs from natural-language prompts through probability distributions of tokens. Various use cases have been proposed; for example, LLMs provide performance comparable to that of veteran data analysts [cheng2023-IsGPTaGoodDataAnalyst] and can be used as LLM-as-a-judge services for data evaluation [kim2023prometheus]. We focus on GPTScore [fu2023gptscore] and G-Eval [liu2023g-eval]. A common problem for LLMs is hallucinations [huang2025hallucination].
LLMs have been utilized in many cybersecurity research, such as data collection [BAYER2023multi-levelfine-tuning, shibli2024AbuseGPT], attack and threat analysis tasks [gadyatskaya2023ChatGPTKnowsYourAttacks, BOFFA2024Logprecis, shibli2024AbuseGPT], and report generation tasks [Gupta2023fromChatGPTtoThreatGPT, mitra2024LocalIntel, perrina2023agir, loumachi2025advancing, wickramasekara2024framework]. However, the above works do not provide a detailed discussion for evaluations of analysis reports. We focus on the evaluation of reports in this paper.
@@ 2.0.2 Security Operation Centers
SOCs address various aspects of cyberattacks, including data collection, determining the presence of intrusions, and responding to threats [Tilbury2024HumansandAutomation]. The current SOCs have monitors industrial control systems, such as factories, as well as the conventional IT systems [albasheer2022Cyber-attack, oba2024ScoreAndYouShallFind]. The major role of SOCs is to detect cyberattacks [gupta2019AutomatedEventPrioritization, renners2019AdaptiveandIntelligible], investigate security alerts [kerstenfield2025tier1analyst, yen2013Beehive] , and make judgments on these alerts [kokulu2019MatchedandMismatchedSOCs]. SOCs roughly consist of two processes, i.e., the evaluation process by analysts and the judgment process by managers. To support and connect both processes, analysis reports are significantly important [kokulu2019MatchedandMismatchedSOCs].
Although SOCs in recent years have introduced automation tools [singh2025contextbuddy, turcotte2025automated, shahjee2022IntegratedNetwork, AGYEPONG2023ASystematicMethod] to reduce workload [Tilbury2024HumansandAutomation], it is still heavy [nepal2024burnout, kerstenfield2025tier1analyst, kersten2024securityalert].Analysts in SOCs also need to understand cyberattacks based on fragmented information [VANDERKLEIJ2022DevelopingDecisionSupport] and frequently require source information to determine appropriate responses [happa2021Assessing]. Although machine learning, such as LLMs, has been developed for analysis in SOCs recently [ANDRADE2019CognitiveSecurity, gonzalez2021SIEM, HUSAK2022Crusoe, mitra2024LocalIntel, albanese2025towardsAI-driven, singh2025contextbuddy], there are limitations for providing analysis reports [michelet2024chatgpt, loumachi2025advancing, wickramasekara2024framework] and security advice [chen2023can] from LLMs.
@@ 2.0.3 Evaluation of Analysis Reports in SOCs
We focus on evaluations of analysis reports in SOCs as the main problem setting. The current report generation tools [loumachi2025advancing, michelet2024chatgpt, wickramasekara2024framework] have limitations due to hallucinations, since LLMs often cause hallucinations and these tools also lack explicit mechanisms for evaluating of analysis reports. Instead, we aim to explore whether LLMs can appropriately evaluate these reports. We note that the evaluation of analysis reports remains challenging: general LLMs may fail to align with human judgment because these reports include domain-specific contexts [chiang2023closer].
@@
We aim to design an evaluation framework that takes analysis reports as input and returns evaluation results as output. The output includes both quantitative scores and qualitative feedback from the perspectives of SOC practitioners, such as veteran analysts. We consider that readers of analysis reports are both analysts and managers, while the generation of the reports is outside the scope of this paper.
@@ 3 Designing Analyst-wise Checklist for Security Report Evaluation
This section aims to clarify the criteria for evaluating the quality of security alert analysis reports and to construct the Analyst-wise Checklist as a quality evaluation metric reflecting these criteria. The following section constitutes the Response to RQ1. This type of report is a highly specialized document that presupposes advanced security domain knowledge. It has been pointed out that general text evaluation metrics are insufficient for accurately assessing the quality of such domain-specific reports [stufflebeam2000guidelines]. The proposed Analyst-wise Checklist integrates evaluation criteria that were iteratively validated through a literature review and interviews, and is designed to focus the assessment on key aspects that directly relate to practical decision-making.
3.1 Construction Process of the Analyst-wise Checklist
The construction of the Analyst-wise Checklist items was performed through the following phased and iterative process. This approach is a common practice for incorporating knowledge base utilization and user needs/field context into design [benton2020usability, kwong2025design, rose2022co]: 1) Collection of Candidate Items: Based on a review of security guidelines and existing related literature, we collected and formulated both the information that should be documented in analysis reports and the initial set of candidate evaluation items. 2) Interviews on Report Quality: Semi-structured interviews were conducted with SOC analysts to extract practical and critical evaluation perspectives on report quality. 3) Development of the Analyst-wise Checklist Draft: We aligned the structural elements of analysis reports extracted from the literature review with the implicit evaluation criteria elicited from SOC analysts through interviews, and, based on these correspondences, newly designed concrete checklist items that are particularly important for assessing analysis report quality. 4) Refinement through Expert Review and Evaluation Tests: For the drafted checklist, we first revised the item structure and wording based on expert review. We then conducted trial evaluations in which multiple authors independently applied the checklist to sample reports, and, by reconciling major rating discrepancies and refining the phrasing of problematic items, confirmed an acceptable level of inter-rater agreement before finalizing the checklist. The details of this checklist are described below.
@@ 3.2 Collection of Candidate Items through Literature Review
The initial phase of developing the checklist items for analysis reports involved a systematic literature review. This review aimed to establish the design foundation of the checklist by ensuring the comprehensiveness and clarity of the information required for reporting alert analysis results. The process consisted of the following steps:
@@
1) Literature Collection: To identify the knowledge required for the checklist items to possess pertinence, official guidelines concerning incident reporting and SOC operations, established by public organizations and expert communities, were set as the collection criteria. A total of 13 documents were targeted, including NIST SP 800-series incident handling guidelines and public information on incident reporting from domestic and international CSIRTs, regulatory bodies, and industry groups [johnson2003handbook, force2020security, johnson2016guide, scarfone2008sp, stouffer2023guide, brenner2007iso, cisa_incident_reporting_2024, enisa_eecc_incident_2021, uscert_incident_guidelines_2015, enisa_incident_management_2010, isogj_soc_textbook_2023, isogj_5w1h_2019, enisa_iot_sm_2019].
@@
2) Extraction of Relevant Content: To ensure efficiency in the extraction process, we combined keyword searching, using terms such as "report," "should be described," "should be included," and "information sharing," with manual reading. This was done to extract sentences that enumerated the information required to be included in a report. For lengthy guidelines spanning dozens to hundreds of pages, we utilized an LLM. Specifically, we queried the LLM with "List the information that should be included in the incident report envisioned by this guideline, and specify where it is written." This provided a candidate list to efficiently narrow down the relevant sections.
@@
3) Manual Selection of Statements: To ensure the clarity and concreteness required for checklist items, the statements obtained from keyword extraction and the LLM candidate lists were manually selected. By cross-referencing with the original text, only descriptions that could genuinely be considered "information items to be included in the report" were retained.
@@
4) Organization of Information: The list of extracted passages was systematically coded by two of the authors with respect to the items that should be documented in an analysis report. Through this coding process, the extracted descriptions were conceptually grouped according to their functional role within the analysis report and organized into three overarching categories and eleven concrete documentation items, thereby establishing the foundation of the checklist. The draft set of three categories and eleven items was then reviewed by two SOC experts and subsequently revised. To ensure the comprehensiveness of the literature base, we extended the initial set of documents with several additional sources. As this extended review did not yield any new content that needed to be extracted, we regarded the data as saturated with respect to the information to be identified from the literature.
We present the set of items that should be documented in an analysis report their categories below.
@@
Decision support and action planning: This category groups report elements that help frontline staff and managers grasp the current situation and determine appropriate next steps. The checklist items assess whether recommended actions and follow-up are concrete, executable, and well-justified, and whether they connect immediate response with longer-term improvement. [enisa_incident_management_2010, isogj_5w1h_2019, isogj_soc_textbook_2023, johnson2003handbook, force2020security, johnson2016guide, scarfone2008sp, uscert_incident_guidelines_2015]. Main report elements: Analysis Status; Impact Assessment; Confirmation Requests; Response Actions; Recurrence Prevention & Lessons.
@@
Accountability and quality assurance of the investigation and analysis process: This category ensures that investigation activities and analytic judgments are documented in a transparent, traceable manner, supporting the reliability and auditability of SOC operations. The checklist items assess whether the process is described systematically and with sufficient detail to enable third-party review, compliance verification, and future incident handling. [stouffer2023guide, enisa_incident_management_2010, cisa_incident_reporting_2024, brenner2007iso, isogj_5w1h_2019, force2020security, johnson2016guide, scarfone2008sp, uscert_incident_guidelines_2015] Main report elements: Investigation & Analysis Methods; Evidence & Supporting Data; Related Policies, Guidelines & Standards.
@@
Technical understanding and root cause clarification of the event: This category covers the technical description of what happened, where, to whom, and how. The checklist items assess whether the incident is described concretely enough to allow technical reconstruction and to justify the reported conclusions. [enisa_incident_management_2010, enisa_iot_sm_2019, enisa_eecc_incident_2021, johnson2003handbook, cisa_incident_reporting_2024, force2020security, scarfone2008sp, uscert_incident_guidelines_2015] Main report elements: Event Description & Interpretation; Root Cause; Incident Source & Impacted Systems; Communication Details; Vulnerability Information.

### v3

@@ 1 Introduction
However, writing analysis reports for security alerts forces a heavy workload on SOC practitioners, because these reports must contain actionable insights [kokulu2019MatchedandMismatchedSOCs]. To reduce this workload, tools, such as large language models (LLMs), that introduce actionable insights into analysis reports are in high demand [albanese2025towardsAI-driven]. Although there are several tools to generate analysis reports for security alerts [loumachi2025advancing, michelet2024chatgpt], automatically generated reports often contain inaccurate information, which may degrade performance in SOCs [kokulu2019MatchedandMismatchedSOCs]. Even when state-of-the-art LLMs are used, they often cause hallucinations, such as inaccurate information [huang2025hallucination]. As another motivation, writing analysis reports to the exact requirements of report-writing guidelines is stressful [nepal2024burnout]. The reason is that evaluation of analysis reports should identify any lack of descriptions as well as judgments from the perspectives of veteran SOC practitioners [RYAN2012Quantifying]. Although the most common mitigation approach for the above stress is to constantly evaluate analysis reports [kokulu2019MatchedandMismatchedSOCs], evaluations of the reports need evaluation criteria based on expert knowledge [kerstenfield2025tier1analyst, kersten2023give]. Moreover, the current guidelines are often abstract and may fail to capture the evaluation criteria from the perspective of veteran SOC practitioners [johnson2016guide, agyepong2020towards]. Namely, even evaluations of analysis reports using LLMs are challenging due to a lack of enough evaluation criteria.
@@
We first design the Analyst-wise Checklist, which contains the evaluation criteria in its specific items to reflect SOC practitioners’ knowledge for evaluating analysis reports in the real world, as the answer to RQ1. To design this, we conducted a literature review of 13 public guidelines and handbooks in the IT and OT domains and semi-structured interviews with 15 SOC practitioners. We also conducted a validation test to confirm the checklist with multiple SOC practitioners. (See Section 3 for details.)
Next, based on the above checklist, we propose a novel framework, called Multi-perspective Evaluation System for Security Analysis using Llm Assistance (MESSALA), to maximize the evaluation of analysis reports in SOCs. In a nutshell, MESSALA guides LLMs in evaluating analysis reports by leveraging the Analyst-wise Checklist from multiple perspectives. As described in Section 4, MESSALA can imitate SOC practitioners’ cognition [zhong2016automate] to obtain expert knowledge from the checklist, and therefore, can return appropriate evaluation results with actionable feedback for each analysis report. When we conduct quantitative evaluations with MESSALA, we show that it can rate analysis reports for the evaluation criteria in comparison with the existing methods [fu2023gptscore, liu2023g-eval]. Thus, to answer RQ2, we confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports from the perspective of veteran SOC practitioners. (See Section 5 for details.)
Third, we conduct qualitative evaluations to identify whether the feedback generated by MESSALA aligns with judgments from the perspective of veteran SOC practitioners. As a result, we confirm if MESSALA can provide more actionable feedback to both novice and veteran SOC practitioners than the existing methods, as it is more specific and easier to understand. It means that MESSALA also enables LLMs to qualitatively evaluate analysis reports through its feedback. (See Section 6 for details.)
@@
  • We design the Analyst-wise Checklist, which contains evaluation criteria for evaluating analysis reports in order to reflect real-world SOC practitioners’ knowledge through literature review and interviews with SOC practitioners, as the answer to RQ1.
  • We propose MESSALA, a novel framework, that maximizes evaluation for each criterion and provides feedback from the perspective of SOC practitioners, and then confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports under the evaluation criteria. This is identical to the answer to RQ2.
@@ 2 Related Work and Our Problem Setting
In this section, we provide background on related work on LLMs and SOCs. We then describe the evaluation of analysis reports in SOCs as the main problem setting.
@@ 2.0.1 Large Language Models
LLMs generate text outputs from natural-language prompts through probability distributions of tokens. Various use cases have been proposed; for example, LLMs provide performance comparable to that of veteran data analysts [cheng2023-IsGPTaGoodDataAnalyst] and can be used as LLM-as-a-judge services for data evaluation [kim2023prometheus]. We focus on GPTScore [fu2023gptscore] and G-Eval [liu2023g-eval]. A common problem for LLMs is hallucination [huang2025hallucination], in which LLMs generate outputs that are inadequate or incorrect.
LLMs have been utilized in many cybersecurity research, such as data collection [BAYER2023multi-levelfine-tuning, shibli2024AbuseGPT], attack and threat analysis tasks [gadyatskaya2023ChatGPTKnowsYourAttacks, BOFFA2024Logprecis, shibli2024AbuseGPT], and report generation tasks [Gupta2023fromChatGPTtoThreatGPT, mitra2024LocalIntel, perrina2023agir, loumachi2025advancing, wickramasekara2024framework]. However, the above works do not provide a detailed discussion of evaluations of analysis reports. We focus on the evaluation of reports in this paper.
@@ 2.0.2 Security Operation Centers
SOCs address various aspects of cyberattacks, including data collection, determining the presence of intrusions, and responding to threats [Tilbury2024HumansandAutomation]. The current SOCs monitor industrial control systems, such as factories, as well as conventional IT systems [albasheer2022Cyber-attack, oba2024ScoreAndYouShallFind]. The major role of SOCs is to detect cyberattacks [gupta2019AutomatedEventPrioritization, renners2019AdaptiveandIntelligible], investigate security alerts [kerstenfield2025tier1analyst, yen2013Beehive] , and make judgments on these alerts [kokulu2019MatchedandMismatchedSOCs]. SOCs roughly consist of two processes, i.e., the evaluation process by analysts and the judgment process by managers. To support and connect both processes, analysis reports are significantly important [kokulu2019MatchedandMismatchedSOCs].
The major challenges of SOCs are to detect novel cyberattacks [gupta2019AutomatedEventPrioritization, renners2019AdaptiveandIntelligible] and reduce the volume of security alerts that analysts must triage [kokulu2019MatchedandMismatchedSOCs, yen2013Beehive, zhong2019LearningfromExpertsExperimence]. Although SOCs in recent years have introduced automation tools [singh2025contextbuddy, turcotte2025automated, shahjee2022IntegratedNetwork, AGYEPONG2023ASystematicMethod] to reduce workload [Tilbury2024HumansandAutomation] for response to security alerts, it is still heavy [nepal2024burnout, kerstenfield2025tier1analyst, kersten2024securityalert].Analysts in SOCs also need to understand cyberattacks based on fragmented information [VANDERKLEIJ2022DevelopingDecisionSupport] and frequently require source information to determine appropriate responses [happa2021Assessing]. Although machine learning, such as LLMs, has been developed for analysis in SOCs recently [ANDRADE2019CognitiveSecurity, gonzalez2021SIEM, HUSAK2022Crusoe, mitra2024LocalIntel, albanese2025towardsAI-driven, singh2025contextbuddy], there are limitations for providing analysis reports [michelet2024chatgpt, loumachi2025advancing, wickramasekara2024framework] and security advice [chen2023can] from LLMs due to several reasons, such as hallucinations described above.
@@ 2.0.3 Evaluation of Analysis Reports in SOCs
We focus on evaluations of analysis reports in SOCs as the main problem setting. The current report generation tools [loumachi2025advancing, michelet2024chatgpt, wickramasekara2024framework] have limitations, because LLMs often cause hallucinations and these tools also lack explicit mechanisms for evaluating of analysis reports. Instead, we aim to explore whether LLMs can appropriately evaluate these reports. We note that evaluating analysis reports remains challenging: general LLMs may fail to align with human judgment because these reports include domain-specific context [chiang2023closer].
@@
We aim to design an evaluation framework that takes analysis reports as input and returns evaluation results as output. The output includes both quantitative scores and qualitative feedback from the perspectives of SOC practitioners, including veteran analysts. We consider that readers of analysis reports are both analysts and managers, while the generation of the reports is outside the scope of this paper.
@@ 3 Designing Analyst-wise Checklist for Security Report Evaluation
This section aims to clarify the criteria for evaluating the quality of security alert analysis reports and to construct the Analyst-wise Checklist that reflects these criteria. The following section constitutes the Response to RQ1. This type of report is a highly specialized document that presupposes advanced knowledge of the security domain. It has been noted that general text evaluation metrics are often insufficient for evaluating the quality of domain-specific reports [stufflebeam2000guidelines]. The Analyst-wise Checklist contains practical evaluation criteria derived from the literature review and interviews and is also effective for decision-making by stakeholders.
3.1 Design Process of the Analyst-wise Checklist
The Analyst-wise Checklist is constructed by the following four processes. This approach is a common practice for incorporating expert knowledge and domain-specific context into design [benton2020usability, kwong2025design, rose2022co]. 1) Collection of Candidate Items: We first collected literature in related research fields and then formulated both the information that should be documented in analysis reports and an initial set of evaluation items. 2) Interviews on Report Quality: Semi-structured interviews were conducted with SOC practitioners to extract practical and critical evaluation criteria with respect to the quality of analysis reports. 3) Development of the Analyst-wise Checklist Draft: Based on the evaluation criteria extracted from the above processes, we newly designed concrete checklist items that are particularly important for evaluating analysis report quality. 4) Refinement through Expert Review and Evaluation Tests: For the checklist draft, we revised its item structure and wording based on expert review, and then conducted trial evaluations in which multiple authors independently applied the checklist to report samples. After refining problematic items, we finalized the checklist items through agreement between the authors and experts. The details of these processes are described below.
@@ 3.2 Collection of Candidate Items through Literature Review
The initial process of designing the checklist involves a literature review. This literature review aims to establish the design foundation of the checklist, and the process consists of the following steps:
@@
1) Literature Collection: To identify the knowledge required for the checklist items, we collected official guidelines for incident response and SOC operations, which were released by public organizations and expert communities. They include NIST SP 800-series incident handling guidelines and public information on incident response from domestic and international CSIRTs, regulatory bodies, and industry groups [meti_is_standards_2016, johnson2003handbook, johnson2016guide, scarfone2008sp, cisa_incident_reporting_2024, enisa_eecc_incident_2021, uscert_incident_guidelines_2015, enisa_incident_management_2010, isogj_soc_textbook_2023, isogj_5w1h_2019, mitre_11strategies_2022, cisa_playbooks_2021, first_csirt_framework_2019].
First, We collected 35 public guidelines and incident handbooks in 2000-2024 from government agencies, leading authorities, and expert communities. These documents were sourced from their official websites and technical repositories. After a screening process based on keyword searches described in Item 2), we selected thirteen representative guidelines.
@@
2) Extraction of Relevant Content: To efficiently extract important checklist items, we combined keyword searching using several words, such as “report," “should be described," “should be included," and “information sharing," with manual reading by the authors. The above step extracts sentences for the information that should be documented in analysis reports. We also use an LLM to follow several guidelines with long pages: specifically, we query the LLM with the prompt as "List the information that should be included in the incident report envisioned by this guideline, and specify where it is written." This provides a candidate set of the checklist items.
@@
3) Manual Selection of Statements: To ensure clear and concrete checklist items, the sentences obtained from the keyword searching and the candidate set of the checklist items are manually checked by the authors themselves. By cross-referencing them with the original texts in each guideline, we selected only the descriptions corresponding to information that should be documented in analysis reports.
@@
4) Organization of Information: The descriptions obtained in the previous step were systematically coded by two of the authors with respect to the items that should be documented in an analysis report. Through this coding process, the extracted descriptions were conceptually grouped according to their functional role within the analysis reports and organized into three categories and eleven concrete items. These three categories and eleven items were reviewed and revised by two SOC practitioners. Afterward, we further collected several guidelines to expand knowledge from the initial collection of literature. Since this expanded step did not yield any new content on the above three categories and eleven items, we regarded the above categories and items as complete with respect to the evaluation criteria. We present these three categories and eleven items that should be documented in analysis reports below.
@@
Decision support and action planning: This category groups report elements that help frontline staff and managers grasp the current situation and determine appropriate next steps. The checklist items assess whether recommended actions and follow-up are concrete and well-justified, and whether they enable immediate response with long-term improvement. In particular, the checklist items include Analysis Status[meti_is_standards_2016, johnson2003handbook, scarfone2008sp, uscert_incident_guidelines_2015, mitre_11strategies_2022, cisa_playbooks_2021, first_csirt_framework_2019], Impact Assessment [johnson2003handbook, johnson2016guide, scarfone2008sp, cisa_incident_reporting_2024, enisa_eecc_incident_2021, uscert_incident_guidelines_2015, enisa_incident_management_2010, isogj_soc_textbook_2023, isogj_5w1h_2019, mitre_11strategies_2022, cisa_playbooks_2021, first_csirt_framework_2019], Confirmation Requests [cisa_incident_reporting_2024, enisa_incident_management_2010, isogj_5w1h_2019, first_csirt_framework_2019], Response Actions [meti_is_standards_2016, johnson2003handbook, scarfone2008sp, cisa_incident_reporting_2024, uscert_incident_guidelines_2015, enisa_incident_management_2010, isogj_soc_textbook_2023, isogj_5w1h_2019, mitre_11strategies_2022, cisa_playbooks_2021, first_csirt_framework_2019], and Recurrence Prevention & Lessons [meti_is_standards_2016, johnson2003handbook, johnson2016guide, scarfone2008sp, cisa_incident_reporting_2024, enisa_eecc_incident_2021, enisa_incident_management_2010, isogj_soc_textbook_2023, isogj_5w1h_2019, mitre_11strategies_2022, cisa_playbooks_2021, first_csirt_framework_2019].
@@
Accountability and quality assurance of the investigation and analysis process: This category ensures that investigation activities and analysis judgments are documented in a transparent and traceable manner for SOC operations. The checklist items assess whether the process is described systematically and with sufficient detail to enable third-party review, compliance verification, and future incident handling. In particular, the checklist items include Investigation & Analysis Methods [johnson2003handbook, scarfone2008sp, enisa_incident_management_2010, isogj_soc_textbook_2023, isogj_5w1h_2019, mitre_11strategies_2022, cisa_playbooks_2021, first_csirt_framework_2019], Evidence & Supporting Data [johnson2003handbook, cisa_incident_reporting_2024, enisa_eecc_incident_2021, uscert_incident_guidelines_2015, isogj_soc_textbook_2023, isogj_5w1h_2019], Related Policies [scarfone2008sp, enisa_eecc_incident_2021, isogj_soc_textbook_2023, isogj_5w1h_2019, cisa_playbooks_2021, first_csirt_framework_2019], and Guidelines & Standards [johnson2003handbook, cisa_incident_reporting_2024, enisa_incident_management_2010, isogj_soc_textbook_2023, first_csirt_framework_2019].
@@
Technical understanding and root cause clarification of the event: This category covers the technical description of what happened, where, to whom, and how. The checklist items assess whether the incident is described concretely enough to allow technical reconstruction and to justify the reported conclusions. In particular, the checklist items include Event Description & Interpretation [johnson2003handbook, johnson2016guide, scarfone2008sp, cisa_incident_reporting_2024, enisa_eecc_incident_2021, uscert_incident_guidelines_2015, isogj_soc_textbook_2023, isogj_5w1h_2019, cisa_playbooks_2021, first_csirt_framework_2019], Root Cause [johnson2003handbook, scarfone2008sp, cisa_incident_reporting_2024, enisa_eecc_incident_2021, isogj_soc_textbook_2023, cisa_playbooks_2021, first_csirt_framework_2019, enisa_incident_management_2010], Incident Source & Impacted Systems [johnson2003handbook, scarfone2008sp, enisa_eecc_incident_2021, uscert_incident_guidelines_2015, isogj_soc_textbook_2023, isogj_5w1h_2019, cisa_playbooks_2021, first_csirt_framework_2019, enisa_incident_management_2010], Communication Details[johnson2016guide, scarfone2008sp, isogj_5w1h_2019], and Vulnerability Information [johnson2016guide, scarfone2008sp, enisa_eecc_incident_2021, enisa_incident_management_2010, isogj_soc_textbook_2023, isogj_5w1h_2019, mitre_11strategies_2022, cisa_playbooks_2021, first_csirt_framework_2019] .